### v2

# Information-Geometric Barycenters for Bayesian Federated Learning

Abstract
Federated learning (FL) is a widely used and impactful distributed optimization framework that achieves consensus through averaging locally trained models. While effective, this approach may not align well with Bayesian inference, where the model space has the structure of a distribution space. Taking an information-geometric perspective, we reinterpret FL aggregation as the problem of finding the barycenter of local posteriors using a prespecified divergence metric, minimizing the average discrepancy across clients. This perspective provides a unifying framework that generalizes many existing methods and offers crisp insights into their theoretical underpinnings. We then propose BA-BFL, an algorithm that retains the convergence properties of Federated Averaging in non-convex settings. In non-independent and identically distributed scenarios, we conduct extensive comparisons with statistical aggregation techniques, showing that BA-BFL achieves performance comparable to state-of-the-art methods while offering a geometric interpretation of the aggregation phase. Additionally, we extend our analysis to Hybrid Bayesian Deep Learning, exploring the impact of Bayesian layers on uncertainty quantification and model calibration.
Index Terms:
Bayesian Federated Learning, Hybrid Bayesian DL, Uncertainty Quantification, Model Aggregation.
Figure 1: The BA-BFL framework.
Deterministic Arithmetic Average Aggregations
FedAvg [1],
FedProx [2], PropFair [3],
GiFair [4], FedLaw [5]
Multiplicative Aggregations of Posteriors
FedPA [6], FOLA [7],
FedEP [8], FedIvon [9]
WB / RKLB
with variance $\to$ 0
RKLB
Figure 2: Mapping of various aggregation methods to their corresponding barycenter formulations.
I Introduction
Federated Learning (FL) has emerged as the de facto standard for decentralized learning, particularly in scenarios requiring strong privacy guarantees. As introduced in [1], an FL system consists of a central server that maintains a global model and interacts with multiple clients, i.e., end-user devices, each holding private local data. FL schemes generally operate in two phases: a local learning phase, where each client trains a local model using its own data, and an aggregation phase, where the locally updated models are sent to the server and combined according to a predetermined rule. These phases repeat iteratively, with the global model from the previous iteration serving as the updated local models in the current iteration.
Aggregation plays a central role in FL, enabling the combination of individual client contributions into a global model while preserving data privacy and ensuring communication efficiency. Although various aggregation strategies have been proposed (e.g., [10]), most known aggregation methods rely on variants of weighted averaging. Notable examples include FedAvg [1] and FedProx [2], which aggregate local models by computing a weighted average of their parameters. The choice of weights is an important design parameter, encoding auxiliary attributes, such as the relative importance of each client’s model, for the overall objective or proportionally to the volume of data available at each client.
A key challenge in FL, and generally in distributed learning systems, is the statistical heterogeneity among participating clients. In real-world scenarios, client datasets rarely satisfy the idealized assumption of independent and identically distributed (i.i.d.) data; instead, they exhibit varying degrees of heterogeneity or distributional shifts across clients. As surveyed in [11, 2], five major forms of heterogeneity are commonly identified: label distribution skew: variations in the representation of specific labels across clients (e.g., under- or over-represented classes); feature distribution skew: differences in the feature distributions associated with the same label; concept drift: the same labels correspond to different features across clients; concept shift: cases where identical samples are labeled differently by different clients, and quantity skew: differences in the number of data samples held by each client. While these heterogeneities are highly relevant in practice, addressing all of them simultaneously remains a major challenge. As a result, most existing FL research targets only a subset of these non-idealities [12, 13, 14]. Given the growing applicability of FL in real-world settings, uncertainty quantification and model calibration are central characteristics for building trustworthy and reliable models. Nonetheless, these aspects remain largely underexplored in the existing research around deterministic FL. A preliminary study on the topic, with specific application to healthcare, is presented in [15], providing an overview of various uncertainty quantification methods within the deterministic FL, later implemented in [16]. It is important to note, however, that the techniques discussed are mainly inspired by Bayesian approaches, including Bayesian ensembles and Monte Carlo dropout.
Improving the reliability of models is where Bayesian learning excels, as Bayesian methods allow for better quantification of the model uncertainty and calibration, rendering it a compelling solution for FL in non-i.i.d. contexts. FedPPD [17] introduces an FL framework that incorporates uncertainty quantification, where each client, in each round, estimates the posterior distribution over its parameters and the posterior predictive distribution (PPD). The PPD is then distilled into a single deep neural network and sent to the server. pFedBayes [18] and Fedpop [19] are also Bayesian approaches with a focus on uncertainty quantification aspects, proposed in the context of personalized Bayesian Federated Learning (BFL). Nonetheless, we emphasize that many of the existing methods [18, 20, 17, 21] rely on variations of weighted averaging of the posteriors’ parameters.
Contributions
This work introduces a unifying perspective on aggregation methodologies within the context of BFL through the lens of barycentric aggregation (BA-BFL). Given a divergence metric, we interpret the aggregation process as a geometric problem, where the global model is identified as the barycenter of the local posteriors. In contrast to existing methods that often rely on heuristic variations of averaging in the parameter space, our approach is theoretically grounded, as it explicitly minimizes the average divergence between the global posterior and the local posteriors. We show that this methodology generalizes several aggregation strategies previously proposed in the literature. Furthermore, we show that the proposed methods preserve the convergence properties of FedAvg, even in non-convex settings (see Theorem III.1). We evaluate our techniques against state-of-the-art Bayesian aggregation methods in terms of accuracy and uncertainty quantification in a heterogeneous setting. The results demonstrate that our approach achieves performance comparable to existing statistical aggregation techniques. To bridge gaps in the BFL literature, and building on insights from Hybrid Bayesian Deep Learning (HBDL) [22], we extend our analysis to examine the impact of limiting the number of Bayesian layers on the performance of various Bayesian aggregation methods.
Notation
Table I summarizes the key symbols and their meanings for clarity and convenience of reference.
Symbol Meaning
$\mathcal{D}_{k}$ Local dataset of client $k$
$\mathcal{D}$ Union of all datasets $\bigcup_{k=1}^{N}\mathcal{D}_{k}$
$D_{\alpha}$ $\alpha$-divergence
$\operatorname*{D_{KL}}$ Kullback–Leibler divergence
$\operatorname*{W_{2}^{2}}2$ Squared 2-Wasserstein distance
$w_{k}$ Weight of client $k$
$f_{k}$ Local objective of client $k$
$f$ Weighted sum $\sum_{k=1}^{N}w_{k}f_{k}$
$\psi_{k}$ Posterior parameters of client $k$’s model
$\psi_{g}$ Posterior parameters of the global model
TABLE I: Summary of notation.
II Background and Related Work
Federated Learning
An FL system [1] consists of a central server and $N$ clients, engaging in an iterative learning process involving server-client communication. For each communication round, the $k^{th}$ client trains its local model, parameterized by $\theta_{k}$, on its private data $\mathcal{D}_{k}$. Subsequently, the model parameters $\theta_{k}$ are sent to the server, which aggregates all the clients’ models to obtain the global model, later sent to the clients to refine their local models for the next communication round. By aggregating the local models of the clients, FL seeks to achieve a global model $\theta^{*}$ on $\mathcal{D}=\bigcup_{k=1}^{N}\mathcal{D}_{k}$, the aggregated dataset from all participating clients.
In general, an FL system considers objective functions of the form
$\min_{{\theta}}f({\theta})=\sum_{k=1}^{N}w_{k}f_{k}({\theta})$ (1)
where $f_{k}({\theta})=\mathbb{E}_{(x,y)\sim\mathcal{D}_{k}}[\mathcal{L}(\theta;(x,y)]$ is the local objective function of the $k^{th}$ client and $w_{k}$ is its associated weight such that $\sum_{k=1}^{N}w_{k}=1$. Minimizing $f_{k}(\theta)$ at each communication round locally results in $\theta_{k}$.
BFL Method Local Bayesian Technique Global Aggregation Method
FedPA [6] MCMC sampling Multiplicative aggregation of posteriors
pFedBayes [18] Variational Inference Weighted averaging of posteriors’ statistics
FVBA [20] Variational Inference Weighted averaging of posteriors’ statistics
FedEP [8] Variational Inference Multiplicative aggregation of posteriors
FedHB [23] Variational Inference Bayesian posterior inference
FOLA [7] Laplace Approximation Multiplicative aggregation of posteriors
Not Named [21] Variational Inference & MC dropout Weighted averaging of posteriors’ statistics
FedPPD [17] MCMC sampling Weighted averaging of posteriors’ statistics
FedIvon [9] IVON Multiplicative aggregation of posteriors
TABLE II: Categorization of parametric Client-Side Bayesian Federated Learning methods.
Hybrid Bayesian Deep Learning (HBDL)
Despite the remarkable results in terms of model performance, deep learning does not address crucial problems in realistic scenarios, such as reliability and uncertainty quantification. In a recent position paper [24], the authors propose Bayesian Deep Learning as a solution for the ethical, privacy, and safety challenges of modern DL. Acknowledging ongoing issues with Bayesian DL, e.g., the added complexity cost of performing Bayesian learning on large-scale deep models, the authors envision the alternative framework of HBDL to maintain the efficiency and lower complexity of deep learning while retaining the reliability of Bayesian DL. HBDL is also discussed in [25] as Bayesian inference on the (n-)last layer(s) only. The idea behind this framework is to substitute part of the layers in a Bayesian deep model with deterministic layers, making them equivalent to the classical deep learning counterpart. The possibility of having only a partially Bayesian model allows for retaining uncertainty quantification capabilities while reducing the complexity compared to a fully Bayesian model.
Bayesian Federated Learning
BFL aims to incorporate the strengths of Bayesian deep learning into the FL framework, offering a potential solution to the challenges outlined earlier. In this work, we focus primarily on parametric client-side BFL methods [26], and we categorize them according to their strategies for global model construction. Table II summarizes this discussion.
  • Multiplicative Aggregation of Posteriors: FedPA [6] employs Stochastic Gradient Markov Chain Monte Carlo (MCMC) for local posterior inference, aggregating client posteriors through a product of Gaussian distributions. FOLA [7] approximates local posteriors using Laplace approximation. A multivariate Gaussian product mechanism is used for global posterior construction, while prior distributions derived from the global posterior guide local training providing a continual learning setting. FedEP [8] frames FL as a distributed Variational Inference problem, aligning the global posterior with local posteriors using multiplicative aggregation. FedIvon [9] adopts Improved Variational Online Newton (IVON) [27] to approximate local posteriors as Gaussians with diagonal covariance, efficiently updating both mean and variance using second-order information. The server aggregates these local posteriors via the weighted product of the local Gaussians. It is important to note that approximating the global posterior as the weighted product of local posteriors is equivalent to computing their geometric mean, which also corresponds to the RKLB aggregation method introduced in this work, offering a geometric explainability to this aggregation’s strategy.
  • Weighted Averaging of Posteriors’ Statistics: pFedBayes [18] employs variational inference to incorporate uncertainty into model parameters. From a continual learning perspective, it minimizes the KL divergence between global and local posterior distributions, balancing local reconstruction error with global alignment, mainly presented in the paper as a personalization aspect. FVBA [20] investigates the aggregation of Variational Bayesian Neural Networks with five statistical aggregation schemes. In [21], the authors integrate Bayesian deep learning with FL, employing variational inference and Monte Carlo Dropout for local models’ inference. As in [20], the different statistical aggregations are evaluated, highlighting the importance of the choice of the aggregation method on the model performance. FedPPD [17] leverages MCMC sampling for local posterior inference and distilling posterior predictive distributions into single deep neural networks using Stochastic Gradient Langevin Dynamics. It adopts either simple averaging or a distillation-based approach for global aggregation.
  • Bayesian Posterior Inference: FedHB [23] introduces a hierarchical Bayesian framework where local model parameters are governed by a global latent variable. variational inference is used to optimize local and global posteriors through a block-coordinate optimization.
III Proposed Method
In this section, we introduce our problem formulation and present the main theoretical results. We start by formalizing the key technical aspects of the client-side BFL framework.
Client-Side BFL
The Bayesian view presents a different framework for FL. The goal is to estimate the posterior distribution of the global model’s parameters, $p(\theta^{*}|\mathcal{D})$, given the posterior distributions of local models $p(\theta_{k}|\mathcal{D}_{k})$.
Nevertheless, exact posterior inference is usually intractable, requiring the use of approximate inference methods instead. In this work, we consider variational inference [28, 29] to approximate the local posteriors given a common prior distribution $p(\theta)$ and the client likelihoods $p(\mathcal{D}_{k}|\theta_{k})$.
For a parametric family $\mathcal{M}_{\Psi}$ parametrized by $\psi\in\Psi$, the optimization problem seeks to identify the distribution $q_{\psi}\in\mathcal{M}_{\Psi}$ that minimizes the KL divergence from the posterior distribution $p(\theta|\mathcal{D})$, i.e.,
$\min_{\psi\in\Psi}\operatorname*{D_{KL}}(q_{\psi}(\theta)\|p(\theta|\mathcal{D}))$ (2)
However, the minimization in (2) is not directly tractable and is commonly approached through the derivation of the Negative Evidence Lower Bound surrogate objective $\min_{\psi\in\Psi}\mathcal{L}(\psi,D)$, where
$\mathcal{L}(\psi,D)=-\mathbb{E}_{q_{\psi}(\theta)}[\log p(\mathcal{D}|\theta)]+\operatorname*{D_{KL}}(q_{\psi}(\theta)||p(\theta)).$ (3)
The local models are trained by minimizing (3) to achieve their models’ posterior distributions $p(\theta_{k}|\mathcal{D}_{k}),~{}\forall k\in\{1,..,N\}$. The local posteriors are then aggregated in order to get the global model’s posterior $p(\theta^{*}|\mathcal{D})$. Given this setting, we now introduce our main assumptions regarding the common prior $p(\theta)$ and the parametric family $\mathcal{M}$, which will stay valid throughout the rest of this paper.
Assumption 1.
For each client, we assume the prior distribution $p(\theta)$ to be a $d$-dimensional Gaussian with independent marginals, parameterized by a zero mean vector $\mathbf{0}_{d}$ and an identity covariance matrix $\mathbf{I}_{d}$.
Assumption 2.
(Mean-field Model) The parametric family $\mathcal{M}_{\Psi}$ is composed of $d$-dimensional Gaussian distributions with independent marginals, i.e., $q_{\psi}=\mathcal{N}(\mu,\Sigma)$, with mean $\mu\in\mathbb{R}^{d}$ and diagonal covariance $\Sigma=\operatorname*{diag}(\sigma^{2}_{1},\ldots,\sigma^{2}_{d})$.
Algorithm 1 BA-BFL: Barycentric Aggregation for Bayesian Federated Learning
0:   Server’s Input: number of communication rounds $R$, aggregation weights $\{w_{i}\}_{i=1}^{N}$, global distribution’s initial parameters $\psi_{g}^{0}$.
0:   Client’s $k$ input: number of local training epochs $T$, local training set $\mathcal{D}_{k}$
0:
1:  for each round $r=1,\dots,R$ do
2:     Sample clients’ subset $\mathcal{S}_{r}\subset\{1,\dots,N\}$
3:     Communicate $\psi_{g}^{r-1}$ to all clients $k\in\mathcal{S}_{r}$
4:     for client $k\in\mathcal{S}_{r}$ do
4:         / $T$ epochs of Gradient Descent (GD) starting at $\psi_{g}^{r-1}$/
5:        $\psi_{k}^{r}\leftarrow\textbf{GD}(\mathcal{L}(\cdot,\mathcal{D}_{k}),\psi_{g}^{r-1}).$      // Eq.3
6:     end for
6:      /Aggregation and global update/
6:     $\psi_{g}^{r}\leftarrow\textbf{D-Barycenter}(\{\psi^{r}_{k}\}_{k=1}^{N},\{w_{k}\}_{k=1}^{N})$
7:  end for
Bayesian Aggregation as Posteriors Barycenter
The main novelty of this work stands in the introduction of the general Barycentric Aggregation framework for BFL (BA-BFL), an aggregation method inspired by the geometric properties of the manifold to which the local posteriors $\{p(\theta_{k}|\mathcal{D}_{k})\}_{k=1\ldots N}$ belong. Given a divergence metric $D$, we propose as a global model the barycenter $p^{*}_{D}$ of the set of clients’ posteriors, i.e., the distribution that minimizes the weighted divergence from a given set. The following problem formalizes this interpretation of the aggregation process.
Problem 1.
($D$-barycenter) Given a statistical manifold $\mathcal{M}$, a divergence function $D:\mathcal{M}\times\mathcal{M}\to[0,\infty)$, and a set of distributions $\mathcal{S}=\{p_{k}\}_{k=1\ldots N}\subseteq\mathcal{M}$ with associated normalized weights $\{w_{k}\}_{k=1\ldots N}$, i.e., $\sum_{k=1}^{N}w_{k}=1$, the barycenter $p^{*}_{D}$ of the set $\mathcal{S}$ is defined as:
$p^{*}_{D}=\operatorname*{arg\,min}_{q\in\mathcal{M}}\sum_{k=1}^{N}w_{k}D(p_{k}||q).$ (4)
We now study Problem 1 under various assumptions on the distribution set $\mathcal{S}$ and the divergence metric $D$. First, we consider the general case where $D=D_{\alpha}$, namely, the divergence belongs to the family of $\alpha$-divergences for $\alpha\in\mathbb{R}\setminus\{0\}$, without any additional assumptions on $\mathcal{S}$. As shown in [30, 31], the corresponding barycenter $p^{*}_{D_{\alpha}}$ takes the following form:
[…]
Moreover, by taking the limit $\alpha\to 0$, i.e., corresponding to the reverse Kullback–Leibler (RKL) divergence $D_{RKL}(p\|q)$, the barycenter $p^{*}_{RKL}$ takes the form:
$p^{*}_{RKL}=\frac{\prod_{k=1}^{N}p_{k}^{w_{k}}}{\int\prod_{k=1}^{N}p_{k}^{w_{k}}d\nu}.$ (6)
We now focus on the case where all $p_{k}\in\mathcal{S}$ are $d$-dimensional Gaussian distributions, i.e., $p_{k}=\operatorname*{\mathcal{N}}(\mu_{k},\Sigma_{k})$, with mean $\mu_{k}$ and covariance matrix $\Sigma_{k}$. This setting derives from Assumption 2, where we assume that the parameters of each Bayesian layer are Gaussian distributed. For the same reasons, we are also interested in the cases where the resulting barycenter is itself Gaussian, to enforce that global and local models belong to the same family of distributions. Alas, this is not the case for the majority of $\alpha$-divergences, as discussed in the following remark.
Remark 1.
(On the $\alpha$-barycenter of a set of Gaussians) Given $\mathcal{S}=\{\operatorname*{\mathcal{N}}(\mu_{k},\Sigma_{k})\}_{k=1\ldots N}$, the barycenter distribution $p^{*}_{D_{\alpha}}$ in (5) is not Gaussian. In fact, $(p^{*}_{D_{\alpha}})^{\alpha}\propto\sum_{k=1}^{N}w_{k}p_{k}^{\alpha}$, showing that the resulting barycenter is related to the Gaussian mixture obtained from the weighted sum of the elements of $\mathcal{S}$. On the other hand, $p^{*}_{RKL}$ is still Gaussian since the Gaussian family is closed under the product operation and (6) is the normalized product of unnormalized Gaussians.
In light of the above technical remark, among the considered $\alpha$-divergences we focus exclusively on the case of $\alpha\to 0$, i.e., $p_{RKL}^{*}$, as the barycenter naturally belongs to the Gaussian family, leaving the study of other $\alpha$-divergences as future work. Given $\mathcal{S}=\{\operatorname*{\mathcal{N}}(\mu_{k},\Sigma_{k})\}_{k=1\ldots N}$, the RKL barycenter is $p_{RKL}^{*}=\operatorname*{\mathcal{N}}(\mu_{RKL},\Sigma_{RKL})$, where
$\Sigma_{RKL}=\left(\sum_{k=1}^{N}w_{k}\Sigma_{k}^{-1}\right)^{-1},~{}\mu_{RKL}=\Sigma_{RKL}\sum_{k=1}^{N}w_{k}\Sigma_{k}^{-1}\mu_{k}.$ (7)
This result is well-established in the literature and has been derived using various approaches (e.g., see [32]).
Similarly to the $D_{RKL}$ divergence, the barycenter of a set of Gaussians in the Wasserstein-2 distance belongs to the same family. In the general setting, the parameters of the barycenter are obtained using a set of fixed-point equations [33]. However, when the set of covariance matrices $\{\Sigma_{k}\}_{k=1\ldots N}$ consists of diagonal matrices, i.e., $\Sigma_{k}=\operatorname*{diag}(\sigma^{2}_{k,1},\ldots,\sigma^{2}_{k,d})$, analytic expressions for the barycenter statistics can be derived, as shown in [33]:
$\Sigma_{\operatorname*{W_{2}^{2}}2}=\left(\sum_{k=1}^{N}w_{k}\Sigma_{k}^{\frac{1}{2}}\right)^{2},\qquad\mu_{\operatorname*{W_{2}^{2}}2}=\sum_{k=1}^{N}w_{k}\mu_{k}.$ (8)
In the sequel, we refer to the aggregation methods resulting from (7) and (8) with the acronyms RKLB and WB, respectively. We discuss the applicability of the proposed methods in HBDL, focusing on cases where part of the architecture is deterministic. In such setting, the posterior distribution $p(\theta_{k,i}|\mathcal{D})$ for the $i^{th}$ layer of the $k^{th}$ client is constrained to be a point-mass located at $\mu_{k,i}$, i.e., $p(\theta_{k,i}|\mathcal{D})=\delta_{(\theta_{k,i}=\mu_{k,i})}$ where $\delta_{x}$ is the Dirac distribution. We investigate the behavior of the proposed methods considering the posterior $p(\theta_{k,i}|\mathcal{D})=\operatorname*{\mathcal{N}}(\mu_{k,i},\epsilon)$ in the limit case of $\epsilon\to 0$. Both (7) and (8) can be shown to be well-defined in the limit, resulting in the barycenter distribution $p^{*}(\theta_{i}|\mathcal{D})=\delta_{\left(\theta_{i}=\sum_{k=1}^{N}w_{k}\mu_{k,i}\right)}$. Notably, this coincides with the arithmetic mean aggregation commonly used in deterministic FL, creating a seamless connection between deterministic and probabilistic aggregation approaches within our framework.
Compared to other state-of-the-art methodologies in parametric client-side BFL - the primary focus of this work - our barycentric perspective extends the widely used weighted multiplication of posteriors, a predominant aggregation method in the literature. We demonstrate that this approach coincides with the RKLB, thereby reinforcing its theoretical foundation. The other baseline used in our comparative studies in Section IV, is the arithmetic mean of the local posteriors’ statistics. In addition, we consider other statistical aggregation methods that, while explored in some comparative studies, have yet to be adopted in practical applications. More broadly, on the server side, BFL can leverage Bayesian ensembles [34], also referred to as Bayesian Model Averaging (BMA), which combines predictions from sampled models to produce a more robust global estimate. Notably, BMA can also be interpreted through the lens of KL barycenters, as highlighted in [Proposition 1.5, [35]]. This barycentric perspective not only provides a strong theoretical grounding for existing methods, as illustrated in Figure 2, but also opens the door to exploring alternative divergence measures to enhance the aggregation process, including the WB.
To conclude this section, we provide theoretical guarantees for the convergence of BA-BFL under both WB or RKLB aggregation, as stated in the following theorem.
Theorem III.1.
(Convergence) Under Assumption 2, and using either RKLB or WB aggregation, BA-BFL inherits and preserves the convergence properties of FedAvg, as shown in [36], for non-convex scenarios with both i.i.d. and non-i.i.d. data.
Proof.
The proof of convergence of BA-BFL is based on existing results on the convergence proof of FedAvg in the non-i.i.d. setup. This connection is possible by recognizing that BA-BFL can be seen as an instance of FedAvg on the parameter space of the chosen parametric family of distributions, subject to a bijective transformation.
Considering a parametric family $\mathcal{M}_{\Psi}$ parametrized by $\psi\in\Psi$, let $F:\Psi\to\hat{\Psi}$ be an invertible mapping. Then, at the $i^{th}$ user, there is no difference between optimizing the local objective on $\Psi$ or a modified version optimized on $\hat{\Psi}$ via the inverse of $F$, i.e.,
$\min_{\psi\in\Psi}f_{i}(\psi)$ $=\min_{\hat{\psi}\in\hat{\Psi}}f_{i}(F^{-1}(\hat{\psi}))$ (9)
$\quad\text{where}\quad f_{i}(\psi)$ $=-\mathbb{E}_{q_{\psi}(\theta)}[\log p(\mathcal{D}|\theta)]+\operatorname*{D_{KL}}(q_{\psi}(\theta)||p(\theta)).$
We are interested in the case where there exists $F$ such that the barycentric aggregation on $\Psi$ induced by a divergence $D$ is equivalent to an arithmetic mean on $\hat{\Psi}$, i.e.,
$\psi_{g}=BA(\{\psi\}_{i=1}^{N})=F^{-1}\left(\sum_{i=1}^{N}w_{i}F(\psi_{i})\right).$ (10)
Under the condition that such mapping exists, then the optimization dynamics of BA-BFL on $\Psi$ is equivalent to FedAvg on $\hat{\Psi}$, i.e., BA-BFL inherits the same convergence properties of FedAvg [^1] For a detailed convergence proof of FedAvg in the non-i.i.d. scenario, we refer the reader to [36].
Considering the family of $d$-dimensional mean-field Gaussian distributions, i.e., with diagonal covariance matrix $\Sigma=\operatorname*{diag}(\{\sigma^{2}_{k}\}_{k=1}^{d})$, it can be parametrized by $\psi=[(\mu_{1},\sigma^{2}_{1}),\ldots,(\mu_{d},\sigma^{2}_{d})]$ with $\mu_{k}\in\mathbb{R}$ and $\sigma_{k}^{2}\in\mathbb{R}^{+}$. Then, we can define the set of entry-wise invertible functions
$F_{RKL}(\mu_{k},\sigma^{2}_{k})=\left(\frac{\mu_{k}}{\sigma^{2}_{k}},\frac{1}{\sigma_{k}^{2}}\right),$ $~{}~{}F^{-1}_{RKL}(\nu_{k},\psi_{k})=F_{RKL}(\nu_{k},\psi_{k})$
$F_{\operatorname*{W_{2}^{2}}2}(\mu_{k},\sigma^{2}_{k})=\left(\mu_{k},\sqrt{\sigma^{2}_{k}}\right),$ $~{}~{}F^{-1}_{\operatorname*{W_{2}^{2}}2}(\nu_{k},\psi_{k})=(\nu_{k},\psi^{2}_{k})$
which satisfy (10) respectively for RKL and $\operatorname*{W_{2}^{2}}2$ divergences. Therefore, given the previous result, BA-BFL under either RKL or $\operatorname*{W_{2}^{2}}2$ divergences enjoys the same convergence properties of FedAvg, thus concluding the proof. ∎
IV Experiments
We devote this section to the experimental investigation of the proposed BA-BFL. To this end, we conduct experimental studies on the FashionMNIST [37], SVHN [38] and CIFAR-10 [39] datasets, within a heterogeneous client setting. The datasets used exhibit varying levels of difficulty.
To compare the proposed methodologies, we consider the following baselines:
  • for deterministic FL, FedAvg [1] aggregates the parameters of the clients’ models through arithmetic weighted average, i.e., $\theta^{*}=\sum_{k=1}^{N}w_{k}\theta_{k}$.
  • for BFL, [20, 21] propose different possible statistical aggregation methods detailed below:
    • Empirical Arithmetic Aggregation (EAA),
      $\mu_{EAA}=\sum_{k=1}^{N}w_{k}\mu_{k},\quad\sigma^{2}_{EAA}=\sum_{k=1}^{N}w_{k}\sigma^{2}_{k}.$
    • Gaussian Arithmetic Aggregation (GAA),
      $\mu_{GAA}=\sum_{k=1}^{N}w_{k}\mu_{k},\quad\sigma^{2}_{EAA}=\sum_{k=1}^{N}w_{k}^{2}\sigma^{2}_{k}.$
    • Arithmetic Aggregation with Log Variance (AALV),
      $\mu_{AALV}=\sum_{k=1}^{N}w_{k}\mu_{k},\quad\sigma_{AALV}^{2}=e^{\sum_{k=1}^{N}w_{k}\log\sigma_{k}^{2}}.$
Metrics
We evaluate the performance of the considered FL algorithms based on three criteria: accuracy, uncertainty quantification, where lower NLL indicates a better model fit, and model calibration, where lower ECE indicates better alignment between predicted probabilities and actual outcomes.
Experimental Setup
To induce label shifts among the $10$ clients participating in the FL scheme, we partition the samples of each label across the clients using a Dirichlet distribution as suggested in [40, 41, 42, 43, 44, 20].
We assign the aggregation weights to reflect the importance of each client in proportion to the volume of data locally owned, i.e., $w_{k}=\tfrac{|\mathcal{D}_{k}|}{|\mathcal{D}|}$ where ||$\cdot$|| ⋅ | indicates the number of samples in the dataset.
The architecture of the global and local models consists of two convolutional layers and three fully connected layers. Following an HBDL approach, we implement the last $n=0,1,2,3$ layers as Bayesian fully connected layers, whereas the remaining layers are deterministic. In our comparative study, increasing $n$ allows measuring the impact of additional Bayesian layers on the uncertainty quantification, model calibration, and the cost-effectiveness of the FL algorithm in time.
TABLE III: Accuracy of the global models resulting from FedAvg, BA-BFL with RKL and $\operatorname*{W_{2}^{2}}2$ barycentric aggregation (RKLB, WB), and BFL baseline aggregation methods (AALV, EAA, GAA). The methods are grouped based on the number of Bayesian layers (Nbl) used in the model architecture. The evaluation is conducted across three datasets (FashionMNIST, CIFAR-10, and SVHN). Bold values represent the highest performance for each dataset, while underlined values denote the best result for the specific dataset within each group of rows corresponding to a specific Nbl.
Nbl Algorithm FashionMNIST SVHN CIFAR-10
0 FedAvg 87.88 $\pm$ 0.79 86.06 $\pm$ 0.45 61.63 $\pm$ 3.11
1 AALV 88.22 $\pm$ 0.34 86.52 $\pm$ 0.29 63.42 $\pm$ 3.02
EAA 88.07 $\pm$ 0.22 86.24 $\pm$ 0.20 63.69 $\pm$ 2.47
GAA 88.15 $\pm$ 0.31 86.36 $\pm$ 0.28 63.66 $\pm$ 2.22
RKLB 88.07 $\pm$ 0.36 86.26 $\pm$ 0.26 63.37 $\pm$ 2.62
WB 88.34 $\pm$ 0.30 86.55 $\pm$ 0.37 63.91 $\pm$ 2.64
2 AALV 87.62 $\pm$ 0.45 85.46 $\pm$ 0.10 65.03 $\pm$ 2.92
EAA 87.53 $\pm$ 0.57 85.64 $\pm$ 0.33 64.02 $\pm$ 1.99
GAA 87.82 $\pm$ 0.64 85.54 $\pm$ 0.44 64.59 $\pm$ 3.51
RKLB 87.59 $\pm$ 0.57 85.57 $\pm$ 0.45 65.20 $\pm$ 3.99
WB 87.69 $\pm$ 0.74 85.57 $\pm$ 0.51 64.74 $\pm$ 3.29
3 AALV 88.07 $\pm$ 0.58 86.15 $\pm$ 0.80 63.71 $\pm$ 3.63
EAA 87.81 $\pm$ 0.54 86.04 $\pm$ 0.62 64.45 $\pm$ 1.79
GAA 88.02 $\pm$ 0.55 86.27 $\pm$ 1.02 64.40 $\pm$ 2.30
RKLB 87.77 $\pm$ 0.80 86.53 $\pm$ 1.03 64.55 $\pm$ 2.97
WB 87.54 $\pm$ 0.54 85.99 $\pm$ 0.68 64.30 $\pm$ 2.55
Overview of the Results
Table 1 summarizes the accuracy performance of BA-BFL using RKLB and WB, alongside the considered baseline methods. All aggregation methods are evaluated under the same model architecture, with an equal number of Bayesian layers. The results show that the proposed aggregation methods achieve performance comparable to the baselines in most scenarios. Furthermore, all Bayesian methods consistently outperform FedAvg across all evaluated datasets. Notably, the best accuracy is obtained with a single Bayesian layer for FashionMINST and SVHN, and with two Bayesian layers for CIFAR-10. Interestingly, increasing the number of Bayesian layers does not always result in improved accuracy.
(a) EAA vs WB
(b) GAA vs WB
(c) RKLB vs WB
(d) AALV vs WB
(e) EAA vs GAA
(f) EAA vs RKLB
(g) GAA vs RKLB
(h) AALV vs EAA
(i) AALV vs GAA
(j) AALV vs RKLB
Figure 3: Bayesian signed-rank test triangle plots comparing aggregation methods on the NLL metric. Each subfigure shows the posterior distribution over the relative performance between two methods.
To assess whether the performance differences between aggregation methods are statistically significant, we employ the Bayesian signed-rank test as described in [45]. For each pair of aggregation methods, we compare their performance across all shared evaluation points, i.e., for each combination of dataset, number of Bayesian layers, and random seed, ensuring a paired analysis under identical experimental conditions. The Bayesian test computes the posterior probabilities that one method outperforms, underperforms, or performs similarly to the other, where similarity is defined within a Region of Practical Equivalence (ROPE). Rather than relying on a fixed ROPE, we adopt a data-driven approach, i.e., for each comparison, the ROPE is set to the 25th percentile of the absolute differences between the two methods. This threshold serves as a conservative estimate of what constitutes a practically negligible performance. When methods behave similarly, most observed differences fall within this region, otherwise, clear performance gaps emerge beyond it. Figure 3 presents the results based on the NLL metric, which is particularly informative as it captures both predictive accuracy and the quality of uncertainty estimation. As shown, the comparisons reveal no statistically significant advantage for any method, indicating broadly comparable performance across all evaluated scenarios.
Figure 4: Effect of Bayesian Layers on Uncertainty Quantification and Model Calibration.
Focusing on the reliability of BFL, we observe that, regardless of the aggregation technique used, the trends reported in Fig. 4 indicate that increasing the number of Bayesian layers in local models improves both global model calibration and uncertainty quantification while reducing ECE and NLL scores. However, this added Bayesian complexity often comes at the expense of reduced time efficiency. As the number of Bayesian layers increases, computational demand rises, leading to longer processing times per communication round. This trade-off between improved model reliability and increased computational cost must be carefully considered in practical applications.
Finally, in Table IV, we compare the performances of BA-BFL against two state-of-the-art parametric client-side BFL methods, pFedBayes [18] and pFedVem [46]. For the proposed method, we consider the setting with RKLB aggregation and three Bayesian layers, as it provides the best performance tradeoff. On the other hand, for pFedBayes and pFedVem, we consider the configurations detailed in their respective original papers. The results show comparable performance across all three methodologies. Our proposed approach achieves relatively better accuracy and NLL on the SVHN and CIFAR-10 datasets. In contrast, pFedVem and pFedBayes yield the best accuracy and NLL, respectively, on FashionMNIST.
Method FashionMNIST SVHN CIFAR-10
pFedVem 89.50 $\pm$ 0.23 86.32 $\pm$ 0.22 60.88 $\pm$ 1.44
pFedBayes 88.02 $\pm$ 0.39 86.03 $\pm$ 0.41 63.86 $\pm$ 1.58
[…]
(a) Comparison in Accuracy
Method FashionMNIST SVHN CIFAR-10
pFedVem 0.45 $\pm$ 0.02 0.80 $\pm$ 0.03 2.44 $\pm$ 0.10
pFedBayes 0.34 $\pm$ 0.02 0.66 $\pm$ 0.04 1.25 $\pm$ 0.06
[…]
(b) Comparison in NLL
TABLE IV: Comparison to state-of-the-art methods
V Conclusions and Future Work
In this paper, we introduced BA-BFL, a novel geometric interpretation of barycenters as a solution to the BFL aggregation problem, providing an explainable aggregation method. The information geometric interpretation of the aggregation step naturally enables operations such as clustering local posteriors directly on the statistical manifold, with potential applications in hierarchical FL. Building on this concept, we recovered two aggregation techniques based on the analytical results for Gaussian barycenters derived using two widely-used divergences: the squared Wasserstein-2 distance and the reverse KL divergence. We demonstrated that BA-BFL retains the same convergence properties as FedAvg for non-convex loss functions, performing robustly across both homogeneous and heterogeneous data scenarios. We experimentally evaluated the proposed methods in a heterogeneous setting, showing gains compared to state-of-the-art methods. Additionally, we examined the impact of varying the number of Bayesian layers in an HBDL context, assessing their effects on accuracy, uncertainty quantification, model calibration, and cost-effectiveness. For future work, we envision several extensions to our study, including expanding the family of distributions to incorporate non-parametric distributions and investigating alternative divergence measures. Furthermore, we plan to address the personalization problem within the framework of barycentric aggregation in BFL.